A tree casts a shadow of 100 inches while a 7 inch branch casts a shadow of 10 inches, how tall is the tree?

Answer by CharlesG2(834)

A tree casts a shadow of 100 inches while a 7 inch branch casts a shadow of 10 inches, how tall is the tree?
shadow 1 / shadow 2 = height 1 / height 2
100 / 7 = x / 10
1000 / 7 = x (multiplied both sides by 10)
142.857 inches approx = x